### v1

@@ 4.2 Convergence analysis for the AMPDA
Theorem 4.9.
Suppose that 2 holds. Then any accumulation point of the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}$, and is a critical point of $F$.
Proof 4.10.
Let $x^{\star}$ be an accumulation point of $\{x^{k}:k\in\mathbb{N}\}$. In view of $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$ (see Proposition 4.5), we derive from the compactness of $\mathcal{X}$ that $x^{\star}\in\mathcal{X}$. Next, we shall show that $x^{\star}$ is a critical point of $F$. Let $\mathcal{K}\subseteq\mathbb{N}$ be a subsequence, such that $\{x^{k}:k\in\mathcal{K}\}$ converges to the $x^{\star}$. Invoking Proposition 4.5 and (33), we have $x^{k+1}\in\mathrm{prox}_{\alpha_{k}c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right)$ for some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. Hence, for any $(x,k)\in\mathcal{C}\times\mathcal{K}$, it holds that
$\alpha_{k}c_{k}f(x^{k+1})+\frac{1}{2}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle\\ \leq\alpha_{k}c_{k}f(x)+\frac{1}{2}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle.$
@@
(52) $c_{k}f(x^{k+1})+\frac{1}{2\overline{\alpha}}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle\\ \leq c_{k}f(x)+\frac{1}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle.$
@@
(53) $c_{\star}f(x^{\star})\leq c_{\star}f(x)+\frac{\|x-x^{\star}\|_{2}^{2}}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}+\left\langle{x-x^{\star}},{\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star}}\right\rangle,$
@@
(54) $x^{\star}\in\arg\min\left\{c_{\star}f_{\mathcal{C}}(x)+\frac{1}{2\alpha}\|x-x^{\star}+\alpha(\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star})\|^{2}_{2}:x\in\mathbb{R}^{n}\right\}$
with $\alpha=\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. By invoking the generalized Fermat’s Rule, the relation (54) yields (28), which together with the fact $x^{\star}\in\mathcal{X}\subseteq\mathrm{dom}(F)$ indicates that $x^{\star}$ is a critical point of $F$ according to Definition 3.9. This completes the proof.
@@
(55) $x^{k+1}\in\arg\min\left\{f(x)+\Big{\langle}\nabla(hg)(x^{k})-\left(\frac{f(x^{k})}{g(x^{k})}+h(x^{k})\right)\nabla g(x^{k}),\right.\\ \left.x-x^{k}\Big{\rangle}+\frac{g(x^{k})}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\right\}.$
@@ 5 Sequential convergence of AMPDA
In this section, we investigate the convergence of the entire sequence generated by the AMPDA for solving problem (1) under the KL assumption. To this end, we first present a key proposition regarding the sufficient descent of $Q$ defined by (31).
Proposition 5.1.
Suppose that 2 holds, and let $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then, there exists $a>0$, such that, for all $k\in\mathbb{N}$,
(56) $Q(x^{k+1},y^{k},z^{k},c_{k+1})+a\|x^{k+1}-x^{k}\|_{2}^{2}\leq Q(x^{k},y^{k-1},z^{k-1},c_{k}).$
Proof 5.2.
Thanks to Proposition 4.5, we know that (34) holds with the $\widehat{x}^{k}=x^{k+1}$ for any $k\in\mathbb{N}$. This, along with (32), implies that (56) holds with $a=\sigma/2$.
Next we prove that the relative error condition stated in Proposition 2.2 (ii) holds for $Q$ at $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$ generated by the AMPDA.
Proposition 5.3.
@@
(57) $\mathrm{dist}(0,\partial Q(x^{k+1},y^{k},z^{k},c_{k+1}))\leq b\|x^{k+1}-x^{k}\|_{2}.$
Proof 5.4.
@@
(58) ^$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})\\ \supseteq\left[\begin{array}[]{l}\left(2\bar{c}+\bar{c}^{2}(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle)\right)\widehat{\partial}f_{\mathcal{C}}(\bar{x})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{y}+\nabla h_{1}(\bar{x})-\bar{z}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{x}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 2f_{\mathcal{C}}(\bar{x})+2\bar{c}f_{\mathcal{C}}(\bar{x})\left(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle\right)\end{array}\right].$
@@
(59) $w_{f_{\mathcal{C}}}^{k+1}:=\frac{g(x^{k})}{\alpha_{k}}(x^{k}-x^{k+1}-\alpha_{k}\nabla h_{1}(x^{k})+\alpha_{k}z^{k})+c_{k}f(x^{k})y^{k}\in\widehat{\partial}f_{\mathcal{C}}(x^{k+1}).$
@@
(60) $w^{k+1}:=(w_{x}^{k+1},w_{y}^{k+1},x^{k}-x^{k+1},w_{c}^{k+1})\in\widehat{\partial}Q\left(x^{k+1},y^{k},z^{k},c_{k+1}\right),$
where $w_{x}^{k+1},w_{y}^{k+1}$ and $w_{c}^{k+1}$ is given as below:
(61) $w_{x}^{k+1}$ $:=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)\right)w_{f_{\mathcal{C}}}^{k+1}$
@@
(62) $w_{y}^{k+1}$ $:=c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})(x^{k}-x^{k+1}),$
(63) $w_{c}^{k+1}$ $:=2c_{k+1}f_{\mathcal{C}}(x^{k+1})\left(g(x^{k+1})+g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle\right).$
@@
(65) $\left\|w_{c}^{k+1}\right\|_{2}\leq 2M_{f/g^{2}}M_{g}(L_{g}+M_{\partial g})\|x^{k+1}-x^{k}\|_{2},$
@@
(66) $w_{x}^{k+1}=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)-c_{k}\right)w_{f_{\mathcal{C}}}^{k+1}\\ +\left(c_{k}^{2}f_{\mathcal{C}}(x^{k})-c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})\right)y^{k}+\nabla h_{1}(x^{k+1})-\nabla h_{1}(x^{k})+\frac{1}{\alpha_{k}}(x^{k}-x^{k+1}).$
@@
(67) $\left|2c_{k+1}+c_{k+1}^{2}\left(g^{*}(y^{k})-\left\langle x^{k+1},y^{k}\right\rangle\right)-c_{k}\right|$
@@
where the first equality follows from $g^{*}(y^{k})=\langle x^{k},y^{k}\rangle-g(x^{k})$. Besides, in view of boundedness of $\left\{\left(x^{k},y^{k},z^{k},c_{k}\right):k\in\mathbb{N}\right\}$ (see Corollary 4.7) and $\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma\leq\alpha_{k}\leq\overline{\alpha}$ (see Proposition 4.5), we deduce from (59) that
@@
$\quad+\left(L_{f/g^{2}}M_{\partial g}+L_{\nabla h_{1}}+(\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma)^{-1}\right)\|x^{k+1}-x^{k}\|_{2},$
@@
for each $k\in\mathbb{N}$, where $L_{f/g^{2}}>0$ and $L_{\nabla h_{1}}>0$ are the Lipschitz moduli of $f/g^{2}$ and $\nabla h_{1}$ on $\mathcal{X}$, respectively. In view of (60), (64), (65) and (68), we conclude that (57) holds for each $k\in\mathbb{N}$. This completes the proof.
Now we are ready to establish the sequential convergence of the AMPDA.
Theorem 5.5.
@@
Suppose that 2 holds. If the $Q$ given by (31) is a proper closed KL function, then the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA converges to a critical point of (1).
Proof 5.6.
Let the sequence $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then, according to Proposition 2.2, Corollary 4.7, Theorem 4.9, Proposition 5.1 and Proposition 5.3, it remains to prove that $Q_{\infty}:=\lim_{k\to\infty}Q(x^{k+1},y^{k},z^{k},c_{k+1})$ exists and $Q\equiv Q_{\infty}$ holds on the set of all accumulation points of $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$. In view of Step 2 of AMPDA and (32), we have $F(x^{k+1})\leq Q(x^{k+1},y^{k},z^{k},c_{k+1})\leq F(x^{k}).$ This together with Corollary 4.7 leads to the existence of $Q_{\infty}$ and $Q_{\infty}=F_{\infty}$.
@@
We let $(x^{\star},y^{\star},z^{\star},c_{\star})$ be an accumulation point of $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$, and $\{(x^{k_{j}+1},y^{k_{j}},z^{k_{j}},c_{k_{j}+1}):j\in\mathbb{N}\}$ be the subsequence converging to $(x^{\star},y^{\star},z^{\star},c_{\star})$. Firstly, we have $c_{\star}=\lim_{j\to\infty}c_{k_{j}+1}=\lim_{j\to\infty}1/g(x^{k_{j}+1})=1/g(x^{\star})$ due to the continuity of $1/g$ on $\mathcal{X}$. Secondly, we know that $x^{k_{j}}\to x^{\star}$ thanks to $x^{k_{j}+1}\to x^{\star}$ and (51). Thirdly, $(y^{\star},z^{\star})\in\partial g(x^{\star})\times\partial h_{2}(x^{\star})$ follows from $(y^{k_{j}},z^{k_{j}})\to(y^{\star},z^{\star})$ with $(y^{k_{j}},z^{k_{j}})\in\partial g(x^{k_{j}})\times\partial h_{2}(x^{k_{j}})$ for each $j\in\mathbb{N}$. Consequently, we have
$Q(x^{\star},y^{\star},z^{\star},c_{\star})\overset{{(I)}}{=}F(x^{\star})\overset{{(II)}}{=}\lim_{j\to\infty}F(x^{k_{j}+1})\overset{{(III)}}{=}F_{\infty},$
@@
where ${(I)}$ follows from (32), ${(II)}$ holds since $F$ is continuous on the compact $\mathcal{X}\subseteq\Omega$, and ${(III)}$ holds thanks to Corollary 4.7. This completes the proof.
Next, we show that problems (2) and (3) meet the requirements on $Q$, and hence the solution sequence generated by AMPDA for the two problems converges to a critical point of them if the initial point $x^{0}$ satisfies (41). First, for problem (2) $g^{*}$ is the indicator function on the $\ell_{2}$ unit ball, while for problem (3) $g^{*}$ is the indicator function on $\{x\in\mathbb{R}^{n}:\|x\|_{\infty}\leq 1,\|x\|_{1}\leq K\}$ (see, for example, [41]). Second, for problems (2) and (3), the corresponding auxiliary objectives $Q$ (see (31)) are in the form of
(69) $Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2c\|x\|_{1}+g^{*}(y)\\ -c^{2}\|x\|_{1}\langle x,y\rangle+\frac{\lambda}{2}\|Ax-b\|_{2}^{2}+h_{2}^{*}(z)-\langle x,z\rangle,$
where $\iota_{\mathcal{C}}$ is the indicator function on $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$, $g^{*}$ is as discussed above, and $h_{2}(x)=\frac{\lambda}{2}\left\|\mathcal{T}_{\mu}(Ax-b)\right\|_{2}^{2}$. Since $h_{2}$ is convex piecewise linear quadratic, $h_{2}^{*}$ is also piecewise linear quadratic ([34, Theorem 11.14]). Invoking this and the equation (69) together with the definition of semi-algebraic functions ([12, Definition 5]), we know that $Q$ is closed and semi-algebraic, and hence $Q$ is a closed KL function ([12, Theorem 3]). In view of the above display and the discussions right after Proposition 4.1, we finally deduce from Theorem 5.5 the following theorem.
Theorem 5.7.

### v2

@@ 4.2 Convergence analysis for the AMPDA
Theorem 4.5
Suppose that Assumption 2 holds. Then any accumulation point of the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}$, and is a critical point of $F$.
Proof
Let $x^{\star}$ be an accumulation point of $\{x^{k}:k\in\mathbb{N}\}$. In view of $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$ (see Proposition 4.3), we derive from the compactness of $\mathcal{X}$ that $x^{\star}\in\mathcal{X}$. Next, we shall show that $x^{\star}$ is a critical point of $F$. Let $\mathcal{K}\subseteq\mathbb{N}$ be a subsequence, such that $\{x^{k}:k\in\mathcal{K}\}$ converges to the $x^{\star}$. In view of Proposition 4.3, (34) and (55), it holds for all $(x,k)\in\mathcal{C}\times\mathcal{K}$ that
$\alpha_{k}c_{k}f(x^{k+1})+\frac{1}{2}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle\\ \leq\alpha_{k}c_{k}f(x)+\frac{1}{2}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})}\right\rangle,$
@@
$c_{k}f(x^{k+1})+\frac{1}{2\overline{\alpha}}\|x^{k+1}-x^{k}\|_{2}^{2}+\left\langle{x^{k+1}-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle\\ \leq c_{k}f(x)+\frac{1}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}\|x-x^{k}\|_{2}^{2}+\left\langle{x-x^{k}},{\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k}}\right\rangle.$ (57)
@@
$c_{\star}f(x^{\star})\leq c_{\star}f(x)+\frac{\|x-x^{\star}\|_{2}^{2}}{2\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma}\\ +\left\langle{x-x^{\star}},{\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star}}\right\rangle,$ (58)
@@
$x^{\star}\in\arg\min\{c_{\star}f_{\mathcal{C}}(x)+\\ \frac{1}{2\alpha}\|x-x^{\star}+\alpha(\nabla h_{1}(x^{\star})-z^{\star}-c_{\star}^{2}f(x^{\star})y^{\star})\|^{2}_{2}:x\in\mathbb{R}^{n}\}$ (59)
with $\alpha=\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. By invoking the generalized Fermat’s Rule, the relation (59) yields (29), which together with the fact $x^{\star}\in\mathcal{X}\subseteq\mathrm{dom}(F)$ indicates that $x^{\star}$ is a critical point of $F$ according to Definition 4. $\square$
@@
$x^{k+1}\in\arg\min\left\{f(x)+\Big\langle\nabla(hg)(x^{k})-\left(\frac{f(x^{k})}{g(x^{k})}+h(x^{k})\right)\nabla g(x^{k}),\right.\\ \left.x-x^{k}\Big\rangle+\frac{g(x^{k})}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\right\}.$ (60)
@@ 5 Sequential convergence of AMPDA
In this section, we investigate the convergence of the entire sequence generated by AMPDA for solving problem (1) under the KL assumption. To this end, we first present the following proposition, which regards the sufficient descent of $Q$ defined by (32), and is a direct result of (33) and the iterative procedure of AMPDA.
Proposition 5.1
Suppose that Assumption 2 holds, and let $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. Then, it holds for all $k\in\mathbb{N}$ that
$Q(x^{k+1},y^{k},z^{k},c_{k+1})+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|_{2}^{2}\leq Q(x^{k},y^{k-1},z^{k-1},c_{k}).$ (61)
Next we prove that the relative error condition stated in Proposition 2.1 (ii) holds for $Q$ at $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$ generated by AMPDA.
Proposition 5.2
@@
$\mathrm{dist}(0,\partial Q(x^{k+1},y^{k},z^{k},c_{k+1}))\leq b\|x^{k+1}-x^{k}\|_{2}.$ (62)
Proof
@@
$\widehat{\partial}Q(\bar{x},\bar{y},\bar{z},\bar{c})\\ \supseteq\left[\begin{array}[]{l}\left(2\bar{c}+\bar{c}^{2}(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle)\right)\widehat{\partial}f_{\mathcal{C}}(\bar{x})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{y}+\nabla h_{1}(\bar{x})-\bar{z}\\ \bar{c}^{2}f_{\mathcal{C}}(\bar{x})\partial g^{*}(\bar{y})-\bar{c}^{2}f_{\mathcal{C}}(\bar{x})\bar{x}\\ \widehat{\partial}h_{2}^{*}(\bar{z})-\bar{x}\\ 2f_{\mathcal{C}}(\bar{x})+2\bar{c}f_{\mathcal{C}}(\bar{x})\left(g^{*}(\bar{y})-\langle\bar{x},\bar{y}\rangle\right)\end{array}\right].$ (63)
@@
$w_{f_{\mathcal{C}}}^{k+1}:=\frac{g(x^{k})}{\alpha_{k}}(x^{k}-x^{k+1}-\alpha_{k}\nabla h_{1}(x^{k})+\alpha_{k}z^{k})+c_{k}f(x^{k})y^{k}\in\widehat{\partial}f_{\mathcal{C}}(x^{k+1}).$ (64)
@@
$w^{k+1}:=(w_{x}^{k+1},w_{y}^{k+1},x^{k}-x^{k+1},w_{c}^{k+1})\in\widehat{\partial}Q\left(x^{k+1},y^{k},z^{k},c_{k+1}\right),$ (65)
where $w_{x}^{k+1},w_{y}^{k+1}$ and $w_{c}^{k+1}$ are given as below:
$w_{x}^{k+1}$ $:=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)\right)w_{f_{\mathcal{C}}}^{k+1}$ (66)
@@
$w_{y}^{k+1}$ $:=c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})(x^{k}-x^{k+1}),$ (67)
$w_{c}^{k+1}$ $:=2c_{k+1}f_{\mathcal{C}}(x^{k+1})\left(g(x^{k+1})+g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle\right).$ (68)
@@
$\left\|w_{c}^{k+1}\right\|_{2}\leq 2M_{f/g^{2}}M_{g}(L_{g}+M_{\partial g})\|x^{k+1}-x^{k}\|_{2},$ (70)
@@
$w_{x}^{k+1}=\left(2c_{k+1}+c_{k+1}^{2}(g^{*}(y^{k})-\langle x^{k+1},y^{k}\rangle)-c_{k}\right)w_{f_{\mathcal{C}}}^{k+1}\\ +\left(c_{k}^{2}f_{\mathcal{C}}(x^{k})-c_{k+1}^{2}f_{\mathcal{C}}(x^{k+1})\right)y^{k}+\nabla h_{1}(x^{k+1})-\nabla h_{1}(x^{k})+\frac{1}{\alpha_{k}}(x^{k}-x^{k+1}).$ (71)
@@
$\left|2c_{k+1}+c_{k+1}^{2}\left(g^{*}(y^{k})-\left\langle x^{k+1},y^{k}\right\rangle\right)-c_{k}\right|$ (72)
@@
where the first equality follows from $g^{*}(y^{k})=\langle x^{k},y^{k}\rangle-g(x^{k})$. In view of boundedness of $\left\{\left(x^{k},y^{k},z^{k},c_{k}\right):k\in\mathbb{N}\right\}$ (see Corollary 4.4) and $\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma\leq\alpha_{k}\leq\overline{\alpha}$ (see Proposition 4.3), we deduce from (64) that
@@
$\|w_{x}^{k+1}\|_{2}\leq M_{w_{f_{\mathcal{C}}}}m_{g}^{-2}(2L_{g}+M_{\partial g})\|x^{k+1}-x^{k}\|_{2}\\ \quad+\left(L_{f/g^{2}}M_{\partial g}+L_{\nabla h_{1}}+(\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma)^{-1}\right)\|x^{k+1}-x^{k}\|_{2},$ (73)
@@
for each $k\in\mathbb{N}$, where $L_{f/g^{2}}>0$ and $L_{\nabla h_{1}}>0$ are the Lipschitz moduli of $f/g^{2}$ and $\nabla h_{1}$ on $\mathcal{X}$, respectively. In view of (65), (69), (70) and (73), we conclude that (62) holds for each $k\in\mathbb{N}$. This completes the proof. $\square$
Now we are ready to establish the sequential convergence of AMPDA.
Theorem 5.3
@@
Suppose that Assumption 2 holds. If the $Q$ given by (32) is a proper closed KL function, then the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA converges to a critical point of (1).
Proof
Let the solution sequence $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ be generated by AMPDA. We prove the convergence of the sequence $\{x^{k}:k\in\mathbb{N}\}$ by verifying all the conditions in Proposition 2.1. Corollary 4.4 (i) shows the boundedness of the solution sequence $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$, and Propositions 5.1 and 5.2 respectively show that the Sufficient Descent Condition and the Relative Error Condition hold for the solution sequence. Thus, it remains only to verify the Continuity Condition.
@@
To verify the Continuity Condition, it suffices to show that the limit $Q_{\infty}:=\lim_{k\to\infty}Q(x^{k+1},y^{k},z^{k},c_{k+1})$ exists and $Q\equiv Q_{\infty}$ holds on the set of all accumulation point of the solution sequence. In view of Step 2 of AMPDA and (33), we have $F(x^{k+1})\leq Q(x^{k+1},y^{k},z^{k},c_{k+1})\leq F(x^{k}).$ This together with Corollary 4.4 (ii) leads to the existence of $Q_{\infty}$ and $Q_{\infty}=F_{\infty}$. We then let $(x^{\star},y^{\star},z^{\star},c_{\star})$ be an accumulation point of $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$, and $\{(x^{k_{j}+1},y^{k_{j}},z^{k_{j}},c_{k_{j}+1}):j\in\mathbb{N}\}$ be the subsequence converging to $(x^{\star},y^{\star},z^{\star},c_{\star})$. Firstly, we have $c_{\star}=\lim_{j\to\infty}c_{k_{j}+1}=\lim_{j\to\infty}1/g(x^{k_{j}+1})=1/g(x^{\star})$ due to the continuity of $1/g$ on $\mathcal{X}$. Secondly, we know that $x^{k_{j}}\to x^{\star}$ thanks to $x^{k_{j}+1}\to x^{\star}$ and (53). Thirdly, $(y^{\star},z^{\star})\in\partial g(x^{\star})\times\partial h_{2}(x^{\star})$ follows from $(y^{k_{j}},z^{k_{j}})\to(y^{\star},z^{\star})$ with $(y^{k_{j}},z^{k_{j}})\in\partial g(x^{k_{j}})\times\partial h_{2}(x^{k_{j}})$ for each $j\in\mathbb{N}$. Consequently, the Continuity Condition follows from
$Q(x^{\star},y^{\star},z^{\star},c_{\star})\overset{{(I)}}{=}F(x^{\star})\overset{{(II)}}{=}\lim_{j\to\infty}F(x^{k_{j}+1})\overset{{(III)}}{=}F_{\infty}=Q_{\infty},$
@@
where (I){(I)} follows from (33), (II){(II)} holds since $F$ is continuous on the compact $\mathcal{X}\subseteq\Omega$, and (III){(III)} follows from Corollary 4.4 (ii).
Combining the above arguments, we conclude that the sequence $\{x^{k}:k\in\mathbb{N}\}$ is convergent. Together with Theorem 4.5, we further conclude that the sequence $\{x^{k}:k\in\mathbb{N}\}$ converges to a critical point of $F$. $\square$
We further note that, if in Definition 1 the function $\phi$ can be chosen as $\phi(z)=a_{0}z^{1-\theta}$ for some $a_{0}>0$ and $\theta\in[0,1)$, then $\varphi$ is said to satisfy the KL property with the exponent $\theta$. It is known from (Bolte-Daniilidis-Lewis-Shiota:2007SIAM-OPT, , Corollary 16) that every proper closed semialgebraic function satisfies the KL property with some exponent $\theta\in[0,1)$. Similar to the analysis in Attouch-Bolte:MP2009 ; Wen-Chen-Pong:2018COA ; Xu-Yin:SIAMIS:13 , the convergence rate of the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA can be estimated by the potential KL exponent of $Q$, and the following statements hold:
  1. (i)
    If $\theta=0$, then the sequence converges in finitely many steps;
  2. (ii)
    If $\theta\in(0,1/2]$, then the sequence converges R-linearly;
  3. (iii)
    If $\theta\in(1/2,1)$, then the convergence is sublinear.
Moreover, by (Rockafellar:70, , Theorem 13.2), we know that, if $g$ is positively homogeneous (i.e., $g(\alpha x)=\alpha g(x)$ holds for all $\alpha>0$ and $x\in\mathbb{R}^{n}$), the conjugate $g^{*}$ is an indicator function of a closed convex set. Using this, for problem (1) we have that $c^{2}f(x)g^{*}(y)=g^{*}(y)$, and the corresponding auxiliary function given by (32) takes the form
$Q(x,y,z,c)=\iota_{\mathcal{C}}(x)+2cf(x)+g^{*}(y)\\ -c^{2}f(x)\langle x,y\rangle+h_{1}(x)+h_{2}^{*}(z)-\langle x,z\rangle.$ (74)
According to the discussions on semialgebraic functions at the end of Section 2.2, we see from (74) that, if the components $f$, $g$, $h_{1}$, $h_{2}$ and the constraint set $\mathcal{C}$ are semialgebraic, then $Q$ is a proper closed semialgebraic function, and hence a KL function. Combining this with the discussions right after Proposition 4.1, we finally invoke Theorem 5.3 to derive the following convergence result for AMPDA when applied to the signal recovery models (2) and (3).
Theorem 5.4